State whether each of the following statements is true or false:
$(1)$ For the probability $P(A)$ of any event $A$,$0 < P(A) < 1$.
$(2)$ The probability of the month of March having $5$ Sundays is $\frac{1}{7}$.
$(3)$ If the probability of an event $A$ occurring is $\frac{3}{7}$,then the probability of event $A$ not occurring is $\frac{4}{7}$.

Vedclass pdf generator app on play store
Vedclass iOS app on app store
(C) $(1)$ False. The probability $P(A)$ of any event $A$ satisfies $0 \le P(A) \le 1$. It can be $0$ (impossible event) or $1$ (sure event).
$(2)$ False. March has $31$ days,which means $4$ weeks and $3$ extra days. The extra days can be (Sun,Mon,Tue),(Mon,Tue,Wed),(Tue,Wed,Thu),(Wed,Thu,Fri),(Thu,Fri,Sat),(Fri,Sat,Sun),or (Sat,Sun,Mon). Out of $7$ possibilities,$3$ contain a Sunday. Thus,the probability is $\frac{3}{7}$.
$(3)$ True. The probability of an event not occurring is $P(\text{not } A) = 1 - P(A)$. Given $P(A) = \frac{3}{7}$,then $P(\text{not } A) = 1 - \frac{3}{7} = \frac{4}{7}$.

Explore More

Similar Questions

In a sample study of $642$ people,it was found that $514$ people have a high school certificate. If a person is selected at random,the probability that the person has a high school certificate is:

$A$ box contains $100$ cards marked with numbers $1$ to $100$. If one card is drawn at random from the box,find the probability that the number on the card is: $(1)$ a multiple of $7$,$(2)$ a number smaller than $30$,$(3)$ a prime number.

Difficult
View Solution

Over the past $200$ working days,the number of defective parts produced by a machine is given in the following table:
Number of defective parts $0$ $1$ $2$ $3$ $4$ $5$ $6$ $7$ $8$ $9$ $10$ $11$ $12$ $13$
Days $50$ $32$ $22$ $18$ $12$ $12$ $10$ $10$ $10$ $8$ $6$ $6$ $2$ $2$

Determine the probability that tomorrow's output will have not more than $5$ defective parts.

$A$ bag contains $40$ balls of the same shape and weight. Among the balls, $18$ are red, $12$ are yellow, and the rest are blue. What is the probability that a ball drawn at random from the bag is blue?

In an oratory competition,$30$ students participated,out of which $12$ are boys. Find the probability that the first prize is won by a girl.

Vedclass Products

For Students

Vedclass Test Series

Mock tests in real JEE/NEET style with performance analysis. 5-day free trial.

Start Free Trial
For Teachers

Exam Paper Generator

Generate Set A/B/C/D exam papers from 7.5L+ questions in 2 minutes. 3 chapters free.

Try Free
For Institutes

Online Exam Module

Live online exams with unlimited students, 360° analytics & white-label branding.

See Demo